# TRATES: Trait-Specific Rubric-Assisted Cross-Prompt Essay Scoring

### v1

@@ Abstract
Sohaila Eltanbouly Qatar University se1403101@qu.edu.qa                        Salam Albatarni Qatar University sa1800633@qu.edu.qa                        Tamer Elsayed Qatar University telsayed@qu.edu.qa
@@ 1 Introduction
As Large Language Models (LLMs) represent the recent advancements in NLP, there is a growing trend towards the integration of LLMs in AES Naismith et al. (2023); Do et al. (2024a). The common approach involves providing the LLM with task descriptions, rubrics, and essays for scoring. However, this method has not achieved the desired results, falling short of basic baseline performance Yancey et al. (2023); Mansour et al. (2024). Other studies employed LLMs as conversational models, where scoring is performed on multiple steps Stahl et al. (2024); Lee et al. (2024). While these approaches improve upon basic LLM prompting, they still lag behind baseline performance. This highlights the need for a hybrid approach integrating the advanced text analysis capabilities of LLMs with the previously well-established AES methods.
@@
At its core, TRATES leverages an LLM to generate and extract trait-specific features given a trait rubric in two stages. Initially, a set of questions is generated from the rubric to assess a specific trait. The LLM then answers each question individually, given the essay. The intuition is to streamline the rubric to facilitate the assessment of various aspects within the trait, rather than assessing the trait as a whole. Finally, trait-specific features are combined with writing-quality and prompt-specific features to train a cross-prompt classical regression model that predicts the trait scores of essays from unseen prompts. This approach allows us to handle the scoring of various traits within a unified framework while automatically tailoring rubric-based features for each specific trait, a task that would be difficult to achieve without the capabilities of LLMs.
Our contribution is five-fold: (1) we present TRATES, a novel rubric-assisted framework for trait-based cross-prompt AES. It provides a simple and generic pipeline that can be used to score any trait while generating features specific to the trait, and combines the strengths of powerful LLMs (for feature generation and extraction) with the simplicity of a basic regression model (for scoring), (2) TRATES establishes a new state-of-the-art (SOTA) performance on all traits on a widely-used dataset, (3) we conduct an ablation study to show the significance of each feature category, (4) we assess the generalizability of TRATES over two different datasets, and (5) we publicly release all trait-specific features to enable future research. [^2] Link will be added soon
@@ 2 Related Work
Research on LLMs for AES is expanding but remains limited. Promising results were demonstrated by Do et al. (2024a, b) with fine-tuning T5 for prompt-specific scoring. Aside from fine-tuning, several LLM prompting techniques were applied to AES Mansour et al. (2024). GPT-4 showed improvement with few-shot examples; however, it failed to outperform a simple XGBoost baseline Yancey et al. (2023). Similarly, GPT-3.5, used for scoring and feedback generation, was unsuccessful Han et al. (2023). Moreover, Stahl et al. (2024) explored various LLM prompting strategies for scoring and feedback generation, including impersonation and chain of thought. Recently, Lee et al. (2024) proposed a Multi Trait Specialization framework that engages the LLM in a conversation to score the essay holistically.
@@ 3 Cross-prompt Trait-based AES
In trait-based AES, a prompt $p$ is defined as a tuple ($a_{p}$, $T_{p}$, $E_{p}$), where $a_{p}$ is the prompt task description, $T_{p}$ is a set $\{(t,r_{t})\}$ of traits; each trait $t$ is associated with a rubric $r_{t}$, and $E_{p}$ is a set $\{(e,\{s_{e,t}\})\}$ of essays written for the prompt $p$; each essay $e$ is associated with a score $s_{e,t}$ for each trait $t\in T_{p}$. While, generally, each prompt has its own trait rubrics, those rubrics are usually common across different prompts for specific traits.
@@ 4 TRATES Framework
In essay scoring, the rubric generally serves as a scoring guide that sets out the criteria for different performance levels. For example, a brief 3-level rubric for the organization trait could be: 2 (logical connections between ideas); 1 (weak connections between ideas); and 0 (no organization evident). Human evaluators, typically teachers, use the rubric to assess the student essays, which enhances grading consistency and transparency. Moreover, it provides students with detailed feedback pinpointing areas for improvement. Given the rubric’s importance in manual essay scoring, our primary objective is to develop an automated rubric-based scoring framework that offers the same benefits as manual scoring, focusing on assessing essay traits.
@@ 4.1 Rubric-based Feature Generation
where $Q_{t}=\{q_{1},q_{2},...,q_{n}\}$ is the set of $n$ sub-traits generated from $r_{t}$. The LLM is prompted to formulate the questions to rate the essay’s aspects as high, medium, or low. The rationale is (1) we opt for simple-to-assess questions, (2) we avoided numeric ratings because the generated questions may not always be formulated so that higher ratings imply better quality, and (3) we prioritize clear and easily understandable responses. The LLM-prompt is shown in Figure 2. We note that the same LLM-prompt template is used with any LLM and trait, with adjustments made to include the specific essay type (if any), trait name, and rubric (as is).
@@ 4.2 Trait-specific Feature Extraction
We next use same LLM $\mathcal{M}$ to answer each sub-trait question $q_{i}\in Q_{t}$, given the LLM prompt instructions $I_{ans}$ (illustrated in Figure 3) and essay $e$.
@@ 4.6 Addressing Cross-prompt Challenges
One challenge with using prompts with different rubrics is that each may have a different scoring range. This poses a challenge when training a regression model, as it requires scores to be standardized within a unified range. A direct scaling method is to apply min-max scaling within each prompt, ensuring that all score ranges are mapped to a common scale. However, this ignores grade level differences, where the quality required for a maximum score varies. For example, an 8th-grade essay might earn a maximum score for clear ideas and basic structure, while a 12th-grade essay requires sophisticated analysis and advanced writing for the same score. A min-max scaling would give the same unified score for both. To address that issue, we propose a score-scaling method based on incremental adjustments relative to the highest grade level; scores for a given grade level are scaled within a fixed range, where the minimum score remains constant while the maximum score decreases by one level for each grade below the highest. This ensures the scaled scores are accurately comparable across prompts with different grade levels and rubrics.
@@ 5 Experimental Setup
To test the generalizability of TRATES, we conducted additional experiments over ELLIPSE dataset Crossley et al. (2023) (Table 6), which comprises about 6.5k essays written by English Language Learners for 44 prompts. Each essay is assessed over 6 traits (cohesion (COH), syntax (SYN), vocabulary (VOC), phraseology (PHR), grammar (GRM), and conventions (CNV)) using a standardized rubric with a scoring range [1-5] and increments of 0.5 points.
@@
For ASAP, we used leave-one-prompt-out cross-validation. The hyperparameters of the models are tuned using sequential hyperparameter tuning by optimizing one hyperparameter at a time while keeping others fixed. The hyperparameters and their search space are listed in Table 7. We used early stopping with a patience of 10 epochs, and ReduceLROnPlateau learning-rate scheduler with a patience of 5 epochs and a factor of 0.1. The batch size was fixed to 128 for all the experiments. LLM’s answers (high/medium/low) were mapped to numerical values (3/2/1). We report the performance of our models on the unseen target prompts for each trait and on average. The performance of ProTACT and Li & Ng baselines is reported as in their corresponding studies for ASAP dataset. For ELLIPSE dataset, we trained the ProTACT model using their implementation. [^5] https://github.com/doheejin/ProTACT All experiments were conducted on a machine with an Intel Xeon Processor and NVIDIA RTX A6000 GPU.
@@
Trait-specific features are generated from each unique rubric. Tables 8 and H show the feature count and examples for ASAP, while Table H presents the same for ELLIPSE.
@@
Previous studies typically normalize features within individual prompts, assuming that feature distribution of the unseen prompt is known Do et al. (2023); Li and Ng (2024b). However, this assumption does not align with real-world scenarios. Instead, we aim to develop a generalizable model that can handle unseen prompts of any size. To achieve this, we use the minimum and maximum values from the training dataset for normalization and consistently apply those values to the test data during inference.
@@ 6 Results and Discussion
In this section, we discuss the results of our experiments addressing 4 research questions: RQ1: Are features generated and extracted via LLMs effective? (6.1), RQ2: Can TRATES with trait-specific and generic features improve the performance? (6.2), RQ3: Which feature category holds greater significance? (6.3), and RQ4: How well does TRATES generalize? (6.4).
@@ 6.3 Feature Category Ablation Study (RQ3)
Table 2: Drop in QWK performance of TRATES with Starling on ASAP when excluding one feature category. Bold and underlined values indicate the most and second-most important categories for each trait respectively.
@@
The trait-specific features stand out as the most significant category across all traits except organization. Notably, these are the only features that vary between traits, underscoring their critical role. This highlights the importance of incorporating trait-specific features to effectively capture the unique requirements of each trait. It is also important to note that these are the only features that are generated automatically, whereas all other features required manual feature engineering.
The prompt-specific category comes next by positively contributing to 6 traits. For WC and SF, the negative impact can be attributed to limited diversity of prompts within those traits. Within the generic features, the length-based category positively contributed to 6 traits; it contains a diverse range of features that pertain to various essay aspects, such as words and sentences, which may explain the negative values observed in some traits, as not all features in the category are equally relevant to every trait. Additionally, readability features had the greatest impact on LNG trait, which is expected as it emphasizes grammar and spelling accuracy.
@@ 6.4 Evaluation on ELLIPSE Dataset (RQ4)
To conduct this study, all the steps of TRATES are repeated with Starling (the best-performing LLM on ASAP). We adopted an 11-fold cross-validation approach, where each fold has 4 prompts, setting 1 unseen fold for testing. [^7] Link to folds will be added soon The results are presented in Table 3. We compare TRATES with 4 baselines: ProTACT, LLM-D, LLM-F, and a feature-based model that is trained on the generic and prompt-specific features (GP-F).
@@
7 Conclusion
In this paper, we introduced TRATES, a cross-prompt rubric-based framework utilizing LLMs, traditional features, and classical regression models for scoring essay traits. The framework is designed to reconceptualize the use of LLMs as feature generators and extractors. The results demonstrate the effectiveness of TRATES yielding new SOTA performance on ASAP dataset and setting the first baseline for ELLIPSE dataset. Our findings underscore the positive impact of the added LLM-based trait-specific features on enhancing the accuracy of trait scoring, emphasizing the necessity for novel models tailored to individual traits. In future work, we plan to extend the framework to holistic scoring and explore different prompting techniques for LLM-based trait-specific feature generation. Moreover, studying the usefulness of the generated sub-traits as feedback to the students remains an important direction for future work.
@@ 8 Limitations
Another limitation of this study is that the effectiveness of the generated questions as feedback for students was not assessed. However, the conceptual framework was discussed with experts in psychometrics and education fields, and the potential value of such feedback was acknowledged. We are dedicated to making these questions available to the community for future research. The efficiency of our proposed framework is another aspect that has not been explored in this study.
@@ Appendix A TRATES LLM Prompts
Figure 3: LLM prompt for feature extraction.
Appendix B The Generic Writing-Quality Features
  1. 1.
    Length-based features, such as the number of words and sentences in the essay; these features are considered the most intuitive indicators for writing quality which have been used extensively over the years Mathias and Bhattacharyya (2018); Chen and Li (2023),
  2. 2.
  3. 3.
    Text variations features covering the usage of different part-of-speech tags (POS) and punctuation.
  4. 4.
  5. 5.
Appendix C ASAP Dataset Statistics
The ASAP++ compromise of 12,978 essays written in English in response to 8 prompts. Each of these prompts has a different number of responses and a different score range. There are a total of 10 different traits; each is covered in a subset of the prompts, except the content trait which is covered in all prompts. It’s important to note that the traits of voice and style are only addressed in one prompt, making them unsuitable for cross-prompt experiments. Table 5 presents the statistics and the details of the dataset.
Appendix D ELLIPSE Dataset Statistics
Appendix E LLMs Selection
We selected three LLMs for our experiments:
  1. 1.
  2. 2.
  3. 3.
@@ Appendix E LLMs Selection
Appendix F Zero-shot LLM Scoring
Figure 4: LLM prompt for Zero-shot LLM direct scoring (LLM-D).
Appendix G Hyperparameters

### v2

@@ Abstract
Sohaila Eltanbouly, Salam Albatarni, Tamer Elsayed Computer Science and Engineering Department, Qatar University, Doha, Qatar {se1403101, sa1800633, telsayed}@qu.edu.qa
@@ 1 Introduction
As Large Language Models (LLMs) represent the recent advancements in NLP, there is a growing trend towards the integration of LLMs in AES Naismith et al. (2023); Do et al. (2024). The common approach involves providing the LLM with task descriptions, rubrics, and essays for scoring. However, this method has not achieved the desired results, falling short of basic baseline performance Yancey et al. (2023); Mansour et al. (2024). Other studies employed LLMs as conversational models, where scoring is performed on multiple steps Stahl et al. (2024); Lee et al. (2024). While these approaches improve upon basic LLM prompting, they still lag behind baseline performance. This highlights the need for a hybrid approach integrating the advanced text analysis capabilities of LLMs with the previously well-established AES methods.
@@
At its core, TRATES leverages an LLM to generate and extract trait-specific features given a trait rubric in two stages. Initially, a set of questions is generated from the rubric to assess a specific trait. The LLM then answers each question individually, given the essay. The intuition is to streamline the rubric to facilitate the assessment of various aspects within the trait, rather than assessing the trait as a whole. Finally, trait-specific features are combined with writing-quality and prompt-specific features to train a cross-prompt classical regression model that predicts the trait scores of essays from unseen prompts. This approach allows us to handle the scoring of various traits within a unified framework while automatically tailoring rubric-based features for each specific trait, a task that would be difficult to achieve without the capabilities of LLMs. Our contribution is five-fold:
  1. 1.
    We present TRATES, a novel rubric-assisted framework for trait-based cross-prompt AES that provides a simple and generic pipeline that can be used to score any trait while generating features specific to the trait. It combines the strengths of powerful LLMs (for feature generation and extraction) with the simplicity of a basic regression model (for scoring).
  2. 2.
    TRATES establishes a new state-of-the-art (SOTA) performance on all traits on a widely-used dataset.
  3. 3.
    We conduct an ablation study to show the significance of each feature category.
  4. 4.
    We assess the generalizability of TRATES over two different datasets.
  5. 5.
    We publicly release all trait-specific features to enable future research. [^2] https://github.com/Sohaila-se/TRATES
The remainder of this paper is organized as follows. Section 2 outlines the related work. Section 3 defines the cross-prompt AES problem. Section 4 details our proposed TRATES framework. Section 5 discusses our experimental setup. Section 6 presents the results and offers a comprehensive analysis. Finally, Section 7 concludes with few suggested future work directions.
@@ 2 Related Work
Research on LLMs for AES is expanding but remains limited. Promising results were demonstrated by Do et al. (2024) with fine-tuning T5 for prompt-specific scoring. Aside from fine-tuning, several LLM prompting techniques were applied to AES Mansour et al. (2024). GPT-4 showed improvement with few-shot examples; however, it failed to outperform a simple XGBoost baseline Yancey et al. (2023). Similarly, GPT-3.5, used for scoring and feedback generation, was unsuccessful Han et al. (2024). Moreover, Stahl et al. (2024) explored various LLM prompting strategies for scoring and feedback generation, including impersonation and chain of thought. Recently, Lee et al. (2024) proposed a Multi Trait Specialization framework that engages the LLM in a conversation to score the essay holistically.
@@ 3 Cross-prompt Trait-based AES
In trait-based AES, a prompt $p$ is defined as a tuple ($a_{p}$, $T_{p}$, $E_{p}$), where $a_{p}$ is the prompt task description (a paragraph or so describing the writing task), $T_{p}$ is a set $\{(t,r_{t})\}$ of traits; each trait $t$ (an aspect of student writing, such as organization or sentence fluency) is associated with a rubric $r_{t}$ (a set of criteria used to evaluate the specific trait $t$), and $E_{p}$ is a set $\{(e,\{s_{e,t}\})\}$ of essays written for the prompt $p$; each essay $e$ is associated with a score $s_{e,t}$ for each trait $t\in T_{p}$. While, generally, each prompt has its own trait rubrics, those rubrics are usually common across different prompts for specific traits.
@@ 4 TRATES Framework
In essay scoring, the rubric generally serves as a scoring guide that sets out the criteria for different performance levels. For example, a brief 3-level rubric for the organization trait could be:
Score 2: Organization and connections between ideas are logically sequenced.
Score 1: Organization and connections between ideas are weak.
Score 0: No organization evident.
Human evaluators, typically teachers, use the rubric to assess the student essays, which enhances grading consistency and transparency. Moreover, it provides students with detailed feedback pinpointing areas for improvement. Given the rubric’s importance in manual essay scoring, our primary objective is to develop an automated rubric-based scoring framework that offers the same benefits as manual scoring, focusing on assessing essay traits.
@@ 4.1 Rubric-based Feature Generation
where $Q_{t}=\{q_{1},q_{2},...,q_{n}\}$ is the set of $n$ sub-traits generated from $r_{t}$. The LLM is prompted to formulate the questions to rate the essay’s aspects as high, medium, or low. The rationale is (1) we opt for simple-to-assess questions, (2) we avoid numeric ratings because the generated questions may not always be formulated so that higher ratings imply better quality, and (3) we prioritize clear and easily understandable responses. The LLM-prompt is shown in Figure 2. We note that the same LLM-prompt template is used with any LLM and trait, with adjustments made to include the specific essay type (if any), trait name, and rubric (as is).
@@ 4.2 Trait-specific Feature Extraction
We next use the same LLM $\mathcal{M}$ to answer each sub-trait question $q_{i}\in Q_{t}$, given the LLM prompt instructions $I_{ans}$ (see Appendix A) and essay $e$.
@@ 4.6 Addressing Cross-prompt Challenges
One challenge with using prompts with different rubrics is that each may have a different scoring range. This poses a challenge when training a regression model, as it requires scores to be standardized within a unified range. A direct scaling method is to apply min-max scaling within each prompt, ensuring that all score ranges are mapped to a common scale. However, this ignores grade level differences, where the quality required for a maximum score varies. For example, an 8th-grade essay might earn a maximum score for clear ideas and basic structure, while a 12th-grade essay requires sophisticated analysis and advanced writing for the same score. A min-max scaling would give the same unified score for both. To address this issue, we propose a score-scaling method based on incremental adjustments relative to the highest grade level; scores for a given grade level are scaled within a fixed range, where the minimum score remains constant while the maximum score decreases by one level for each grade below the highest. This ensures the scaled scores are accurately comparable across prompts with different grade levels and rubrics.
@@ 5 Experimental Setup
To test the generalizability of TRATES, we conducted additional experiments over the ELLIPSE dataset Crossley et al. (2023) (Table 7), which comprises about 6.5k essays written by English Language Learners for 44 prompts. Each essay is assessed over 6 traits (cohesion (COH), syntax (SYN), vocabulary (VOC), phraseology (PHR), grammar (GRM), and conventions (CNV)) using a standardized rubric with a scoring range [1-5] and increments of 0.5 points.
@@
For ASAP, we used leave-one-prompt-out cross-validation. The hyperparameters of the models are tuned using sequential hyperparameter tuning, where one hyperparameter is optimized at a time while keeping others fixed. The hyperparameters and their search space are listed in Table 8. We used early stopping with a patience of 10 epochs, and ReduceLROnPlateau learning-rate scheduler with a patience of 5 epochs and a factor of 0.1. The batch size was fixed to 128 for all the experiments. LLM’s answers (high/medium/low) were mapped to numerical values (3/2/1). We report the performance of our models on the unseen target prompts for each trait and on average. The performance of ProTACT and Li & Ng baselines is reported as in their corresponding studies for ASAP dataset. For ELLIPSE dataset, we trained the ProTACT model using their implementation. [^5] https://github.com/doheejin/ProTACT
@@
Trait-specific features are generated from each unique rubric. Tables 9 and A show the feature count and examples for ASAP, while Table A presents the same for ELLIPSE.
@@
Previous studies typically normalize features within individual prompts, assuming that the feature distribution of the unseen prompt is known Do et al. (2023); Li and Ng (2024b). However, this assumption does not align with real-world scenarios. Instead, we aim to develop a generalizable model that can handle unseen prompts of any size. To achieve this, we use the minimum and maximum values from the training dataset for normalization and consistently apply those values to the test data during inference.
@@ 6 Results and Discussion
In this section, we discuss the results of our experiments addressing 4 research questions: RQ1: Are features generated and extracted via LLMs effective? (6.1), RQ2: Can TRATES with trait-specific and generic features improve the performance? (6.2), RQ3: Which feature category holds greater significance? (6.3), and RQ4: How well does TRATES generalize? (6.4). Furthermore, we discuss the inference cost in Section 6.5.
@@ 6.3 Feature Category Ablation Study (RQ3)
Table 2: Drop in QWK performance of TRATES with Starling on ASAP when excluding one feature category. Bold and underlined values indicate the most and second-most important categories for each trait, respectively.
@@
The trait-specific features stand out as the most significant category across all traits except organization. Notably, these are the only features that vary between traits, underscoring their critical role. This highlights the importance of incorporating trait-specific features to effectively capture the unique requirements of each trait. It is also important to note that these are the only features that are generated automatically, whereas all other features require manual feature engineering.
The prompt-specific category comes next by positively contributing to 6 traits. For WC and SF, the negative impact can be attributed to the limited diversity of prompts within those traits. Within the generic features, the length-based category positively contributed to 6 traits; it contains a diverse range of features that pertain to various essay aspects, such as words and sentences, which may explain the negative values observed in some traits, as not all features in the category are equally relevant to every trait. Additionally, readability features had the greatest impact on LNG trait, which is expected as it emphasizes grammar and spelling accuracy.
@@ 6.4 Evaluation on ELLIPSE Dataset (RQ4)
To conduct this study, all the steps of TRATES are repeated with Starling (the best-performing LLM on ASAP). We adopted an 11-fold cross-validation approach, where each fold has 4 prompts, setting 1 unseen fold for testing. [^7] https://github.com/Sohaila-se/TRATES The results are presented in Table 3. We compare TRATES with 4 baselines: ProTACT (a variant of ProTACT without the topic coherence (TC) features, as their extraction was not included in the provided implementation), LLM-D, LLM-F, and a feature-based model that is trained on the generic and prompt-specific features (GP-F).
@@
6.5 Inference Cost
ORG WC SF PA NAR LNG CNV CNT
Num. of Trait-specific Features 25 9 13 18 20 17 15 31
Essay Average Length 350 350 350 100 100 100 350 350
Trait-specific FE Time (msec) 5,655 2,017 2,933 2,252 2,477 2,084 3,379 7,004
Generic + Prompt FE Time (msec) 139 139 139 43 43 43 139 139
Regression Model Time (msec) 0.12 0.12 0.12 0.12 0.12 0.12 0.12 0.12
Total Time (msec) 5794 2156 3072 2295 2520 2127 3518 7143
Table 4: Average inference time of TRATES, using the Starling LLM, per essay of prompt P1 for the ORG, WC, SF, CNV, and CNT traits, and P3 for the other traits. ’FE’ denotes Feature Extraction stages. Time is in milliseconds.
In addition to TRATES predictive performance, the practical deployment of TRATES depends on its inference efficiency. We evaluate the inference cost of TRATES in terms of processing time, focusing on the feature extraction stages and essay scoring.
Table 4 presents the average inference time per essay for each trait with the Starling LLM. For the ORG, WC, SF, CNV, and CNT traits, we report the average inference time over the 1,783 essays of prompt P1. For PA, NAR, and LNG traits, the reported times are averaged over the 1,726 essays of prompt P3. All timings were obtained on an Azure VM equipped with an NVIDIA A10 GPU and an AMD EPYC 74F3 24-Core Processor. The model was loaded using the Hugging Face Transformers library with FP16 precision.
As expected, the majority of the inference cost is attributed to LLM-based feature extraction, with times ranging from 2.02 to 7 seconds, depending on the number of features per trait. Traits from P3 show lower inference times compared to those from P1, primarily due to the shorter essay length, which reduces the LLM prompt size and speeds up processing. In contrast, the time for extracting the generic and prompt-specific features is 0.04 seconds for P3 and 0.14 seconds for P1, whereas the final scoring step using the neural network regression model is extremely fast and consistent across all traits, taking only 0.12 milliseconds. These results highlight that the LLM component highly dominates the overall inference cost.
These findings demonstrate the feasibility of deploying TRATES in applications with moderate-latency requirements. Although latency remains a limitation and warrants further research, AES is not considered a real-time task and can tolerate the current processing delays.
7 Conclusion and Future Work
In this paper, we introduced TRATES, a cross-prompt rubric-based framework utilizing LLMs, traditional features, and classical regression models for scoring essay traits. The framework is designed to reconceptualize the use of LLMs as feature generators and extractors. The results demonstrate the effectiveness of TRATES yielding new SOTA performance on ASAP dataset and setting the first baseline for ELLIPSE dataset. Our findings underscore the positive impact of the added LLM-based trait-specific features on enhancing the accuracy of trait scoring, emphasizing the necessity for novel models tailored to individual traits.
In future work, we plan to extend the framework to holistic scoring and explore different prompting techniques for LLM-based trait-specific feature generation. Moreover, studying the usefulness of the generated sub-traits as feedback to the students remains an important direction for future work.
@@ 8 Limitations
Another limitation of this study is that the effectiveness of the generated questions as feedback for students was not assessed. However, the conceptual framework was discussed with experts in psychometrics and education fields, and the potential value of such feedback was acknowledged. We are dedicated to making these questions available to the community for future research. Moreover, the efficiency remains a limitation of our proposed framework and is an important area for future improvement.